How many moles of each reactant are needed to produce 2.5 mol of aluminum oxide by the following reaction? 4 Al + 3 O2 → 2 Al2O3
1
Expert's answer
2020-07-07T14:34:09-0400

According to the reaction, n(O2) = 3/2 x n(Al2O3) = 3/2 x 2.5 = 3.8 mol;

n(Al) = 2 x n(Al2O3) = 2 x 2.5 = 5 mol.

Answer: 5 moles of Al and 3.8 moles of O2.